OpenStudy (minecraftqueen):

A. 3 B. 4 C. 9 D. 81 3log3(-5) = A. -5 B. 3 C. 9 D. undefined If f(x) = -2x2 + 8x - 4, which of the follwoing is true? A. The maximum value of f(x) is -4. B. The graph of f opens upward. C. The graph of f has no x-intercept D. f is not a one to one function. If f(x) = 5 - 2x, then f-1(-3) = A. 8/39 B. -3 C. 3 D. 32 If logx(3) = 1/4, then x = A. 81 B. 1/81 C. 3 D. 9

OpenStudy (minecraftqueen):

If f(x) = -x2 + 1, then f(x + 1) = A. -x2 + 1 B. -x2 - 2x C. -x2 D. -x2 -2x - 2 If f(x) = x - 4, then (fof)(3) = = A. 1 B. -1 C. -5 D. 5 If ln(3x - 2) = 1, then x = A. 2/3 B. (2 + e)/3 C. 3/2 D. e/3 The number of sol